# Convergence Rate of the Join-the-Shortest-Queue System

### v3

@@ 1 Introduction
Efficient load balancing is central to optimizing resource allocation and minimizing job delays in modern service systems and data centers. The Join–the-Shortest-Queue (JSQ) policy is a popular load balancing algorithm due to its simplicity and optimality under various criteria. Since its introduction by Winston [20], it has received significant attention over the decades. However, the analysis of the JSQ system remains highly nontrivial [5]. Due to the complexity of the problem, most prior works consider a JSQ system only in steady state. Further, researchers typically conduct analysis in some asymptotic regimes, such as the heavy-traffic regime. For example, there are various bounds for the queue length of a JSQ system in its steady state, such as [15, Theorem 10.2.3]. However, traffic patterns in modern data centers tend to be time-varying [2]. This necessitates careful finite-time analysis of JSQ performance, rather than relying only on steady-state behavior. Motivated by this, we focus on analyzing the transient behavior of a JSQ system. As a starting point, we consider a system consisting of two servers with both interarrival and service times following exponential distributions. We analyze its non-asymptotic convergence rate toward its steady state.
As discussed, prior works focus on analyzing the JSQ policy under the heavy-traffic regime and in its steady state. However, unsurprisingly, simulation results in Figure 1 reveal that the convergence rate of a JSQ system can be quite slow when the traffic intensity $\rho$ approaches 1. Thus, to conduct more careful analysis of the JSQ systems, we need to understand their transient behavior. In addition, prior results are not applicable to light-traffic settings since they rely on state space collapse (SSC), which is only associated with the heavy-traffic regime. In this paper, we address this challenge and obtain an explicit upper bound on the difference between the transient distribution of a JSQ system and its steady state for any traffic intensity $\rho\in(0,1)$.
@@
We focus on analyzing the distribution $\mathbb{P}^{t}_{\bm{x}}$ of the queue length vector at time $t$, for a JSQ system with initial queue length $\bm{x}$. In particular, we consider its total variation (TV) distance to the steady state $\pi$ at every time $t\geq 0$. The total variation distance is a common metric quantifying the convergence rate of a stochastic system and is known to be closely connected to many other metrics. For example, bounds on TV distances imply bounds on Hellinger distances. In addition to the TV distance, we derive the convergence rate of the mean queue length. Using coupling, the key challenge in our method is to obtain an upper bound on the expected hitting times or first passage times of the JSQ system. For general continuous-time Markov chains (CTMCs), analyzing the hitting time usually requires extensive exploitation of the structure of the CTMC [4]. We use several structures of the JSQ system, and show that the problem can be reduced to providing an upper bound on $\mathbb{E}[T_{0}]$, where $T_{0}$, to be defined in (3.1), is the time it takes for a JSQ system’s queue length vector to hit the origin starting from state $(1,0)$. After upper bounding $\mathbb{E}[T_{0}]$, we use Markov’s inequality to arrive at a bound on $d_{\mathrm{TV}}\left(\mathbb{P}^{t}_{\bm{x}},\pi\right)$.
@@ 1.2 Related work
Since the pioneering work of Winston [20], the JSQ policy has been extensively studied in the literature [3, 5]. Transient analyses are generally complicated, even for simple systems such as $M/M/1$ queues [16, 1] and $M/M/n$ queues [8]. There are numerical results for transient analysis of JSQ systems [9], but few theoretical results are known regarding their convergence rates. In this work, we analyze the transient behavior through a hitting time analysis (see, e.g., [10]). Though there are some results related to hitting times or first passage times for JSQ systems [21, 17], they cannot directly be used for our problem. For a simple system such as an $M/M/1$ system, one can obtain the moment generating function (MGF) of the hitting time for the queue length to hit state 0 using a construction of an exponential martingale and the optional stopping theorem [14, Proposition 5.4]. One of the key reasons this works is the special property of an $M/M/1$ queue: starting from state $x\geq 1$, the queue length must first hit states $x-1,x-2,\cdots,1$ before hitting state 0. However, this property is not true for a JSQ system. For example, consider a JSQ system with an initial queue length $(x,y)$ with $x,y\geq 1$. It is possible that the queue length vector never hits state (x1,y1)(x-1,y-1) before hitting $(0,0)$. This is one of the technical barriers to analyzing a JSQ system. One possible approach is to use the exponential martingale constructed in [17]; however, that work focuses on the hitting time to $(K,K)$, $K\geq 1$, whereas our interest is in the hitting time to $(0,0)$, which is fundamentally different, making those results inapplicable.
@@ 2.2 Main result
Consider a JSQ system as described in Theorem 1 with $\rho=\frac{\lambda}{2\mu}<1$ and initial queue length vector $\bm{x}$. Then
@@ 3.3 Proof of Lemma 1
We consider a JSQ system with queue length vector $\bm{s}$, driven by the same arrival and potential service processes as $\bm{p}$ and $\bm{q}$ with $s_{j}(0)=\max\left\{q_{j}(0),p_{j}(0)\right\},j\in\{1,2\}$. We couple the three systems $\bm{p},\bm{q},\bm{s}$ similarly to Section 3.1. Formally, let $\bm{s}$ evolve according to
@@
Since $\bm{s}(0)\geq\bm{p}(0)$ and $\bm{s}(0)\geq\bm{q}(0)$ componentwise, Lemma 2 (applied to the coupled pairs $(\bm{s},\bm{p})$ and $(\bm{s},\bm{q})$) implies that for all $t\geq 0$, $\bm{s}(t)\geq\bm{p}(t)$ and $\bm{s}(t)\geq\bm{q}(t)$ componentwise. Hence, for any $t\geq 0$, $\bm{s}(t)=(0,0)$ implies $\bm{p}(t)=(0,0)$ and $\bm{q}(t)=(0,0)$. From Lemma 3, the time $\bm{s}$ hits 0, denoted by $T^{\prime}$, is stochastically dominated by $\sum_{i=1}^{s_{1}(0)+s_{2}(0)}T_{0i}$, where $T_{0i}$ are i.i.d. copies of $T_{0}$ (3.1). Therefore,
@@
Using $\max\left\{x,y\right\}\leq x+y$ for $x,y\geq 0$, (3.5) implies
@@ 3.4 Proof of Lemma 2
      $dp_{2}(t_{1})=1,dq_{1}(t_{1})=1.$
      Combining the above equation and the definition of the JSQ policy, we arrive at $p_{1}(0)>p_{2}(0)$ and $q_{2}(0)\geq q_{1}(0)$. Combining this and the assumption $p_{2}(0)\geq q_{2}(0)$, we have
@@
      $dp_{1}(t_{1})=1,dq_{2}(t_{1})=1.$
      Combining the above equation, $p_{1}(0)\geq q_{1}(0)$, and the definition of the JSQ policy, we arrive at
@@ 3.5 Proof of Lemma 3
Consider two coupled JSQ systems with queue length processes $\bm{p}(t)$ and $\bm{q}(t)$ as governed by (3.2). Suppose $\bm{p}(0)=(x+1,y)$ and $\bm{q}(0)=(x,y)$ for some $x\geq 0,y\geq 0$. Then from Lemma 2, for any $t\geq$0, $\bm{q}(t)=(0,0)$ implies $\bm{p}(t)\in\left\{(1,0),(0,1),(0,0)\right\}$.
Next, we show $T(0,1){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T(1,0){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T_{0}$. We can couple two copies of queues $\bm{q}^{(1)}$ and $\bm{q}^{(2)}$ started from $(1,0)$ and $(0,1)$ by using the same arrival process and the same potential service process for the unique initial job (so that if the job completes first, it completes in queue 1 for $\bm{q}^{(1)}$ and in queue 2 for $\bm{q}^{(2)}$ at the same time). Under this coupling, at the first event time, the two copies will be in the same post-jump state, either both at $(0,0)$ or both at $(1,1)$. Thus, $T(0,1){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T(1,0){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T_{0}$.
@@ 3.6 Proof of Lemma 4
$q^{\text{J}}_{i}(t)\leq q^{\text{M}}_{i}(t),i\in\{1,2\},t\geq 0\text{ almost surely}.$
@@
Next, consider a two-queue JSQ system $\bm{q}^{\text{J}}(t)=(q_{1}(t),q_{2}(t))$ (with an arrival rate $\lambda$ and each queue has a service rate $\mu$) and the product chain $\bm{q}^{\text{M}}(t)=(q^{\text{M}}_{1}(t),q^{\text{M}}_{2}(t))$ of two independent $M/M/1$ queues defined previously, started from the same initial state $\bm{q}^{\text{J}}(0)=\bm{q}^{\text{M}}(0)=(1,0)$. By Proposition 4, there exists a copuling such that

### v4

@@ 1 Introduction
Efficient load balancing is central to optimizing resource allocation and minimizing job delays in modern service systems and data centers. The Join–the-Shortest-Queue (JSQ) policy is a popular load balancing algorithm due to its simplicity and strong performance under various criteria. Since its introduction by Winston [20], it has received significant attention over the decades. However, the analysis of the JSQ system remains highly nontrivial [5]. Due to the complexity of the problem, most prior works consider a JSQ system only in steady state. Further, researchers typically conduct analysis in some asymptotic regimes, such as the heavy-traffic regime. For example, there are various bounds on the queue length of a JSQ system in its steady state, such as [15, Theorem 10.2.3]. However, traffic patterns in modern data centers tend to be time-varying [2]. This necessitates a careful finite-time analysis of JSQ performance, rather than relying only on steady-state behavior. Motivated by this, we focus on analyzing the transient behavior of a JSQ system. As a starting point, we consider a system consisting of two servers with both interarrival and service times following exponential distributions. We analyze its non-asymptotic convergence rate toward its steady state.
As discussed, prior works focus on analyzing the JSQ policy under the heavy-traffic regime and in its steady state. However, unsurprisingly, simulation results in Figure 1 reveal that the convergence rate of a JSQ system can be quite slow when the traffic intensity $\rho$ approaches 1. Thus, to conduct more careful analysis of the JSQ systems, we need to understand their transient behavior.Many prior results rely on state space collapse phenomena in heavy traffic and do not directly yield explicit finite-time bounds for light traffic. In this paper, we address this challenge and obtain an explicit upper bound on the difference between the transient distribution of a JSQ system and its steady state for any traffic intensity $\rho\in(0,1)$.
@@
We focus on analyzing the distribution $\mathbb{P}^{t}_{\bm{x}}$ of the queue length vector at time $t$, for a JSQ system with an initial queue length vector $\bm{x}$. In particular, we consider its total variation (TV) distance to the steady state $\pi$ at every time $t\geq 0$. The total variation distance is a common metric quantifying the convergence rate of a stochastic system and is known to be closely connected to many other metrics. For example, bounds on TV distances imply bounds on Hellinger distances. In addition to the TV distance, we derive the convergence rate of the mean queue length. Using coupling, the key challenge in our method is to obtain an upper bound on the expected hitting times or first passage times of the JSQ system. For general continuous-time Markov chains (CTMCs), analyzing the hitting time usually requires extensive exploitation of the structure of the CTMC [4]. We use several structural properties of the JSQ system, and show that the problem can be reduced to providing an upper bound on $\mathbb{E}[T_{0}]$, where $T_{0}$, to be defined in (3.1), is the time it takes for a JSQ system’s queue length vector to hit the origin starting from state $(1,0)$.
@@ 1.2 Related work
Since the pioneering work of Winston [20], the JSQ policy has been extensively studied in the literature [3, 5]. Transient analyses are generally complicated, even for simple systems such as $M/M/1$ queues [16, 1] and $M/M/n$ queues [8]. There are numerical results for transient analysis of JSQ systems [9], but few theoretical results are known regarding their convergence rates. In this work, we analyze the transient behavior through a hitting time analysis (see, e.g., [10]). Though there are some results related to hitting times or first passage times for JSQ systems [21, 17], they cannot be used directly for our problem. For a simple system such as an $M/M/1$ system, one can obtain the moment generating function (MGF) of the hitting time for the queue length to hit state 0 using a construction of an exponential martingale and the optional stopping theorem [14, Proposition 5.4]. One of the key reasons this works is the special property of an $M/M/1$ queue: starting from state $x\geq 1$, the queue length must first hit states $x-1,x-2,\cdots,1$ before hitting state 0. However, this property is not true for a JSQ system. For example, consider a JSQ system with an initial queue length $(x,y)$, where $x,y\geq 1$. It is possible that the queue length vector never hits state (x1,y1)(x-1,y-1) before hitting $(0,0)$. This is one of the technical barriers to analyzing a JSQ system. One possible approach is to use the exponential martingale constructed in [17]; however, that work focuses on the hitting time to $(K,K)$, $K\geq 1$, whereas our interest is in the hitting time to $(0,0)$, which is fundamentally different, making those results inapplicable.
@@ 2.2 Main result
Consider a JSQ system as described in Theorem 1 with $\rho=\frac{\lambda}{2\mu}<1$ and an initial queue length vector $\bm{x}$. Then
@@ 3.3 Proof of Lemma 1
We consider a JSQ system with queue length vector $\bm{s}$, driven by the same arrival and potential service processes as $\bm{p}$ and $\bm{q}$, with $s_{j}(0)=\max\left\{q_{j}(0),p_{j}(0)\right\},j\in\{1,2\}$. We couple the three systems $\bm{p},\bm{q},\bm{s}$ similarly to Section 3.1. Formally, let $\bm{s}$ evolve according to
@@
From Lemma 3, the time $\bm{s}$ hits 0, denoted by $T^{\prime}$, is stochastically dominated by $\sum_{i=1}^{s_{1}(0)+s_{2}(0)}T_{0i}$, where $T_{0i}$ are i.i.d. copies of $T_{0}$ (3.1). Since $\bm{s}(0)\geq\bm{p}(0)$ and $\bm{s}(0)\geq\bm{q}(0)$ componentwise, Lemma 2 (applied to the coupled pairs $(\bm{s},\bm{p})$ and $(\bm{s},\bm{q})$) implies that for all $t\geq 0$, $\bm{s}(t)\geq\bm{p}(t)$ and $\bm{s}(t)\geq\bm{q}(t)$ componentwise. Thus, for any $t\geq 0$, $\bm{s}(t)=(0,0)$ implies $\bm{p}(t)=(0,0)$ and $\bm{q}(t)=(0,0)$, or $T\leq T^{\prime}$ almost surely. Since $T^{\prime}\preceq\sum_{i=1}^{s_{1}(0)+s_{2}(0)}T_{0i}$, we have
@@
Using $\max\left\{x,y\right\}\leq x+y$ for $x,y\geq 0$ and $\bm{q}(0){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}\pi$, (3.5) implies
@@ 3.4 Proof of Lemma 2
      $dp_{2}(t_{1})=1,dq_{1}(t_{1})=1.$ (3.8)
      Combining (3.8) and the definition of the JSQ policy, we arrive at $p_{1}(0)>p_{2}(0)$ and $q_{2}(0)\geq q_{1}(0)$. Combining this and the assumption $p_{2}(0)\geq q_{2}(0)$, we have
@@
      $dp_{1}(t_{1})=1,dq_{2}(t_{1})=1.$ (3.10)
      Combining (3.10), $p_{1}(0)\geq q_{1}(0)$, and the definition of the JSQ policy, we arrive at
@@ 3.5 Proof of Lemma 3
Consider two coupled JSQ systems with queue length processes $\bm{p}(t)$ and $\bm{q}(t)$ as governed by (3.2). Suppose $\bm{p}(0)=(x+1,y)$ and $\bm{q}(0)=(x,y)$ for some $x\geq 0,y\geq 0$. From Lemma 2, for any $t\geq 0$, $\bm{q}(t)=(0,0)$ implies $\bm{p}(t)\in\left\{(1,0),(0,1),(0,0)\right\}$.
Next, we show that $T(0,1){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T(1,0){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T_{0}$. We can couple two copies of queues $\bm{q}^{(1)}$ and $\bm{q}^{(2)}$ started from $(1,0)$ and $(0,1)$ by using the same arrival process and the same potential service processes for the unique initial job (so that if the job completes first, it completes in queue 1 for $\bm{q}^{(1)}$ and in queue 2 for $\bm{q}^{(2)}$ at the same time). Under this coupling, at the first event time, the two copies will be in the same post-jump state, either both at $(0,0)$ or both at $(1,1)$. Thus, $T(0,1){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T(1,0){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T_{0}$.
@@ 3.6 Proof of Lemma 4
$\sum_{i=1}^{2}q^{\text{J}}_{i}(t)\leq\sum_{i=1}^{2}q^{\text{M}}_{i}(t),t\geq 0\text{ almost surely}.$
@@
Next, consider a two-queue JSQ system $\bm{q}^{\text{J}}(t)=(q^{\text{J}}_{1}(t),q^{\text{J}}_{2}(t))$ (with an arrival rate $\lambda$ and each queue has a service rate $\mu$) and the product chain $\bm{q}^{\text{M}}(t)=(q^{\text{M}}_{1}(t),q^{\text{M}}_{2}(t))$ of two independent $M/M/1$ queues defined previously, started from the same initial state $\bm{q}^{\text{J}}(0)=\bm{q}^{\text{M}}(0)=(1,0)$. By Proposition 4, there exists a coupling such that